OpenStudy (anonymous):

Help please: Bob has two savings accounts. He deposited $200 more into account A than account B. After a period of time, account A has earned $84 in interest at 3%, and account B has earned $24 in interest at a rate of 2%. Write an equation to represent the situation. Explain each variable used. How much money did Bob initially deposit into each account? Solve the equation

OpenStudy (anonymous):

With steps please

OpenStudy (anonymous):

@radar Could you help please

OpenStudy (anonymous):

Can you help? @camerondoherty

OpenStudy (radar):

This could be done several different ways, I prefer to do it using only one variable. Let x = amount of dollars deposited in Account B x + 200 = amount of dollars deposited in account A From information provided by the problem we can now write: .03(x + 200) = $84 Solve for x .03x + 6 = $84 .03x = $78 x=$2,600.00 (account B) x+$200 = $2600 + $200 = $2800 (account A)

OpenStudy (radar):

@ Snippy1470 You tagged me and didn't stick around to follow the steps or ask questions. Shame on you.

OpenStudy (anonymous):

Sorry @radar I have been checking for the last hour. Thanks anyways.